OpenStudy (keke):

simpify; 3 over √(2)

OpenStudy (anonymous):

3V2/2

OpenStudy (anonymous):

\[3\sqrt{2}/2\]

OpenStudy (anonymous):

what is with the word 'simplify'? someone should smack these math teacher. if they mean "write in simplest radical form" then a) they should say it and b) they should say what exactly it means. what is "simpler" about \[\frac{3\sqrt{2}}{2}\]?